Two sides of a triangle are formed by the vectors A = 3i + 6j - 2k and B = 41- j + 3k. Determine the angles of the triangle.

Grade:12th pass

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student, 
We know that in vector,  area of the parallelogram formed by A & B is the magnitude of their cross product. So, area of required triangle is half the magnitude of cross product of A & B. 
If we find A×B , using determinant: 
|i +j +k| 
|3 +6 -2| 
|4 -1 +3| 
=16i-17j-27k. 
Magnitude of A×B =sqrt(16^2 + 17^2 + 27^2) = sqrt(1274) 
So magnitude of area of triangle = sqrt (1274)/2 =17.85 sq units [Ans.]
Hope this helps.
Thank you, 
Soumya Ranjan Mohanty.